OpenStudy (anonymous):

cot(45-A).cot(45-B)=1

OpenStudy (primeralph):

To do what?

OpenStudy (anonymous):

prove.. this question

OpenStudy (amriju):

prove it...or find solutions.?? if solutions r needed..thn we have cot (45-A).cot(45-B)=1...so..cot(45-A)=tan(45-B)...right...now remember tan A=cot(90-A)..so we rewrite this as tan(90-(45-A))=tan(45-B)...so we have..tan(45+A)=tan(45-B)...so you can find the relation betn A nd B..using general solutions..

OpenStudy (amriju):

trigo gen solns u'll get in any standrd maths book..

OpenStudy (anonymous):

how cot(45-A)=tan(45-B)??

OpenStudy (anonymous):

be fast!!

OpenStudy (amriju):

remember..1/tan x=cot x nd vice versa...so cot(45-A).cot(45-B)=1...that implies that cot(45-A)=(1/cot(45-B))...thats cot(45-A)=tan(45-b)..getting it..??

OpenStudy (anonymous):

nopes.. :-(

OpenStudy (amriju):

u do get this that cot(45-A)=(1/cot(45-B))...don't u..??

OpenStudy (anonymous):

got it... i proved it... thnks!!

OpenStudy (amriju):

:)

OpenStudy (anonymous):

now next question!!! i had posted .... r8 now

OpenStudy (amriju):

i got no ques...in the forum...can u pste it here..??

OpenStudy (amriju):

btw which class r u in now??

OpenStudy (anonymous):

11th...

OpenStudy (amriju):

okkkk...so ur ques..paste it here...plz..

OpenStudy (anonymous):

cot(45-A).cot(45+A)=1..prove it

OpenStudy (amriju):

okkk...so use the same property here...cot(45-A)=tan(90-(45-A))...lyk i told u in the last ques..so cot(45-A)=tan(45+A)...so our expression becomes..tan(45+A)*cot(45+A)...so lyk I told u...cotx*tanx is always 1...so proved..

OpenStudy (amriju):

got it??

OpenStudy (anonymous):

yup... now next.. :-) if u dnt mind then...

OpenStudy (amriju):

no but i deserve a medal..:P :P

OpenStudy (anonymous):

i didnt know how to give someone medal.. in this site!! :-p

OpenStudy (amriju):

lol..no probs...type ur ques here..but...along with typing in the forum..in case i fail..

OpenStudy (anonymous):

okk..

OpenStudy (anonymous):

cot2x.cotx-cot2xcot3x-cot3xcotx=1

OpenStudy (amriju):

okk..m on it...ths will tke tym..

OpenStudy (anonymous):

hmm.....

OpenStudy (amriju):

cot(a+b) = (cosa cosb - sina sinb) / (sina cosb + sinb cosa) =cosa cosb / (sina cosb + sinb cosa) - sina sinb / (sina cosb + sinb cosa) = 1/ (sina/cosa + sinb/cosb) - 1/(cosb/sinb + cosa/sina) = 1/ (tana + tanb) - 1/ (cota + cotb) So cot(3x) = 1/(tanx + tan2x) - 1/(cotx + cot2x) so cotx.cot2x - cot2x.cot3x - cot3x.cotx =cotx.cot2x - (cot2x+cotx).cot3x =cotx.cot2x - (cot2x+cotx).(1/(tanx + tan2x) - 1/(cotx + cot2x)) =cotx.cot2x - (cotx+cot2x)/(tanx+tan2x) + 1 Put m = cotx, and n=cot2x = mn - (m+n)/(1/m + 1/n) + 1 = mn - (m+n)/(n/mn + m/mn)) + 1 = mn - (m+n)/((m+n)/mn) + 1 = mn - mn + 1 = 1 get it...??...ws a long one..i used some formulae that can be derieved from simple ones..

OpenStudy (anonymous):

wait let me try!!

OpenStudy (amriju):

okkk...

OpenStudy (anonymous):

got it/... next!!!

OpenStudy (amriju):

yeah okay...hey u preparing for iit??...whr r u frm??

OpenStudy (anonymous):

\[\tan2\theta+ \tan3\theta-\tan5\theta=-\tan2\theta.\tan3\theta.\tan5\]

OpenStudy (amriju):

5theta rite?..not tan 5..

OpenStudy (anonymous):

yup... by mistake.. i had wrtten

OpenStudy (anonymous):

tell me in simply way.. plzzz

OpenStudy (amriju):

this will probabbly use tan(a+b+c)...so i'm not so sure it'll be very simple...although m working on it..nd m nt sure..

OpenStudy (anonymous):

nopes its wriitn that the formila cot(A+B)or cot(A-B)...

OpenStudy (amriju):

yeah my mistake...look tan 5A=tan(2A+3A)....remember the formula for tan(x+y)?..tan(x+y)=(tan x +tan y)/(1-tanx.tany)...so now..tan 5A=tan(2A+3A)..thts ( tan 2A + tan 3A)/(1-tan 2A.tan 3A) thn tan 5A(1-tan 3A tan 2A) = tan 3A +tan 2A tan 5A -tan 5A tan 3A tan 2A = tan 3A +tan 2A tan 5A - tan 3A - tan 2A = tan 5A.tan 3A tan 2A..got it??...sory i wnt to the toilet..so lt rply

OpenStudy (anonymous):

let me try!! :)

OpenStudy (amriju):

:)yeah...hey but r u indian..??

OpenStudy (anonymous):

yup.... ad u??

OpenStudy (amriju):

me too...from whr r u..???

OpenStudy (anonymous):

next ... cotA.cotB+cotB.cotC+cotC.cotA=1

OpenStudy (anonymous):

if A+B+C= pie...

OpenStudy (amriju):

yes..tht was d conditn...

OpenStudy (amriju):

ok...A+B=180-C...rite??...so..tan(A+B)=tan(180-C)...now tan(180-C)=-tan C remember...??...so (tan A + tan B)/(1-tan A tan B) = - tan C or tan A + tan B + tan C = tan A tan B tan C divide both sides by tan A tan B tan C...this is very common process in mny sums.. so 1/tan B 1/ tan C + 1/tan A 1/ tan C + 1/ tan A 1/tan B = 1 or cot B cot C+ cot A cot C + cot A cot B = 1

OpenStudy (amriju):

got it??

OpenStudy (anonymous):

yup.....

OpenStudy (anonymous):

tan18+tan27 tan18tan27=1

OpenStudy (amriju):

should dere be a+ u missd..chck the ques..

OpenStudy (anonymous):

nopes...

OpenStudy (amriju):

tan 27.tan 18. tan27 shud be (tan27)^2.tan18...rite??

OpenStudy (anonymous):

not got....

OpenStudy (amriju):

didn't get me..??? u wrote tan18+tan 27tan 18tan 27=1...so tan 27 tan 18tan27 shud be (tan 27)^2tan18..rite???...did u mean dat...?

OpenStudy (anonymous):

tan18+tan27 +tan18tan27=1

OpenStudy (anonymous):

the quest is this...

OpenStudy (amriju):

yupppp....now it can be done..

OpenStudy (amriju):

so...tan 18+tan27=1-tan 18tan 27..rite...so...(tan 18+tan 27)/(1-tan 18.tan 27)=1...but see...dats the formula for tan(x+y)...lyk (tan x+tan y)/(1-tan xtan y)...correct??...so u getn tan (27+18)=1...thats tan 45=1...which is true offcourse ...got it??

OpenStudy (anonymous):

yup....

OpenStudy (amriju):

so next prob?

OpenStudy (anonymous):

\[\sqrt{3}(\tan140-\tan110)= 1+\tan 140.\tan110\]

OpenStudy (amriju):

hey...this one is easy...lyk the last one... just rearrange it... (tan 140-tan 110)/(1+tan 140.tan 110)=(1/(3)^1/3) thats the form of tan (x-y) remember...?...so tan (140-110)=(1/(3)^1/3)...thats tan 30=(1/(3)^1/3)...thats true...

OpenStudy (anonymous):

wats this????

OpenStudy (amriju):

the proof...??u didn't get it??

OpenStudy (anonymous):

na

OpenStudy (amriju):

ok..so 1.44(tan 140-tan 110)=1+tan 140.tan 110...so we can write (tan 140-tan 110)/(1+tan 140tan 110)=1/1.44=.69...see the left hand side is of the form tan (x-y) thats (tan x-tan y)/(1+tan x.tan y)...so we can write tan (140-110)=.69...thats tan 30=.69...which is true...u get it now...Btw 1.44 is cube root of 3...

OpenStudy (amriju):

hey u got this??

OpenStudy (anonymous):

no.. leave it...

OpenStudy (amriju):

how come u r not getting this...its easy blv me...tan(x-y)=(tan x-tan y)/(1+tanx.tany)...u have to rearrange the terms to get this form...see if u can get that...it wud be betr...if i could actually show u on pen nd paper..